Suppose monthly market demand: P=$50+0.1(I)-0.01Q Where I is consumer’s monthly disposable income. Calculate the quantity demanded given that the current price in the market is $10 and the monthly disposable income is $1,500? Given your answer from (a) and given that I=$1,500, what should be the change in P if you want to increase the demand by 10%?
Answer #1

Answer.a)

If P=10,I=1500

10=50+150-0.01Q

0.01Q=190=19000

b)new demand=19000+10% of 19000=20900.

P=50+150-209

P=-9, which is negative.Price cannot go negative(from 10 to -9, decrease is -19).

Hence even at price=0, demand cannot increase by 10%.